The base of a triangle is 2x+4 and the area is 2x^2+5x+2 sq. meters. How much longer is the base than the height?
Answer by nerdybill(7384)

The base of a triangle is 2x+4 and the area is 2x^2+5x+2 sq. meters. How much longer is the base than the height?
.
since
(1/2)bh = area
(1/2)(2x+4)h = 2x^2+5x+2
(x+2)h = 2x^2+5x+2
h = (2x^2+5x+2)/(x+2)
h = ((2x+1)(x+2))/(x+2)
h = 2x+1
.
How much longer is the base than the height?
base-height
(2x+4)-(2x+1)
2x+4-2x+1
4+1
5 meters